OpenStudy (anonymous):

Solve the open sentence. –2 ≤ n + 4 ≤ 7

OpenStudy (anonymous):

A. n ≥ 6 and n ≤ –4 B. n ≤ 2 and n ≤ 7 C. n ≤ –6 and n ≤ 4 D. n ≥ –6 and n ≤ 3

OpenStudy (sidsiddhartha):

well u have –2 ≤ n + 4 ≤ 7 divide it into two parts -2<=n+4 so from here \[-6\le n\] and n+4<=7 from here \[n \le 3\]

OpenStudy (sidsiddhartha):

got that? @eflatt13

OpenStudy (anonymous):

Solve the open sentence. 8 < 2b OR 2b + 15 < 5 A. b > 4 OR b < –5 B. b < –2 OR b > 3 C. b > –1 OR b < 3 D. b > –3 OR b < 4

OpenStudy (sidsiddhartha):

one question at a time buddy first close it then ask separately :)